what will be lewis dot structure of N2and O2+? and how we can find thier bond order?

Grade:12

1 Answers

Vikas TU
6 years ago
N have 8 valence electrons                                           02 have 14  valence electrons
lewis dot structure would be:                                            lewis dot structure would be:
[N – N]                                                                                                 [O – O]2-
Bond order = [(Number of electrons in bonding molecules) - (Number of electrons in antibonding molecules)]/2   
             For N2+ = ( 10 -4 ) /2 = 3                                                                                               For O2 = (12 -7)/2 = 2.5